OpenStudy (moneybaby980):

will medal+fan An oil company fills 1 over 12 of a tank in 1 over 4 hour with gasoline. At this rate, which expression can be used to determine how long will it take for the tank to fill completely? 1 over 4 ⋅ 1 over 12 1 over 12 ⋅ 4 1 over 4 ⋅ 12 12 ⋅ 4

OpenStudy (anonymous):

If it takes .25 hours to fill 1/12, then how long to fill up 12/12 of the tank. So we write...

OpenStudy (anonymous):

(1/4) / (1/12) = (x)/(12/12)

OpenStudy (moneybaby980):

ok

OpenStudy (anonymous):

x is the amount of hours to fill the barrel given the ratio of the barrel filling up per unit time. So solve for x.

OpenStudy (moneybaby980):

A

OpenStudy (anonymous):

Show me how you got that.

OpenStudy (moneybaby980):

I just thought if u multiply 1/4 times 1/12 u will get the answer

OpenStudy (anonymous):



OpenStudy (anonymous):

How can I rewrite this ?

OpenStudy (moneybaby980):

put 1/4 ⋅1/12= x ???

OpenStudy (anonymous):

When you divide by a fraction, you multiply by its reciprocal. For example

OpenStudy (anonymous):